# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v5

@@ 3 Bridge-Building and Intersectional Liberation
Liberalism’s greatest insight is that individual freedom matters. Anarcho-accelerationism shares this commitment without reservation. Where we diverge is on the conditions for genuine autonomy. A person who depends on a monopolistic employer for income, a monopolistic landlord for shelter, and a monopolistic platform for communication is not free in any meaningful sense (Pettit, 1997). Genuine autonomy requires material independence, which in turn requires commons-based infrastructure: community-owned solar microgrids (Farrell, 2016), self-hosted federated services that eliminate dependence on corporate platforms (Zulli et al., 2020), and platform cooperatives that eliminate dependence on extractive intermediaries (Scholz, 2016).
@@
Intersectionality, as articulated by Crenshaw (1989) and developed by Collins (2019), identifies that systems of oppression do not operate independently but intersect: race, gender, class, sexuality, disability, age, nationality, and other axes of identity produce compound disadvantages that cannot be understood by examining any single dimension in isolation. We argue that UDR, implemented through the decentralized robotic commons, constitutes the most comprehensive intersectional intervention yet proposed, because it eliminates the material basis of all axes of oppression simultaneously. UDR eliminates class by definition: when every person has direct access to goods and services from the robotic commons, the distinction between those who own the means of production and those who must sell their labor dissolves. Racial oppression loses its material leverage when employment, housing, food, healthcare, and education are provided by the commons independently of market participation. The gendered division of labor loses its material foundation when all instrumental labor is automated. When access to resources is decoupled from productive capacity, the distinction between able and disabled bodies loses its economic significance. Economic dependence on discriminatory employers is eliminated for LGBTQ+ individuals. Both the young and the elderly gain full access to the resources of the commons. And when material well-being is no longer tied to citizenship in a wealthy state, the economic desperation that drives forced migration is eliminated. Table˜6 summarizes these interventions.
@@
This progressive path embodies what we term synthetic liberty, in contrast to the negative liberty of neoliberalism. Isaiah Berlin’s classic distinction (Berlin, 1969) between negative liberty (freedom from interference) and positive liberty (freedom to act) illuminates the inadequacy of the neoliberal conception of freedom. Van Parijs (1995) argues that real freedom requires the material conditions for genuine self-determination. Synthetic liberty is the positive freedom that emerges when commons-based infrastructure provides those conditions. It is synthetic in two senses: it is constructed through deliberate institutional and technological design, and it synthesizes the insights of multiple traditions.
@@ 4 The Liberation Stack, Accelerationist Ecologism, and the Pareto Frontier
Anarcho-accelerationism proposes a concrete layered architecture for sociotechnical systems that we term the Liberation Stack, illustrated in Figure˜7. Each layer is composed of existing, mature, open-source technologies.
@@
No community is truly autonomous if it depends on external entities for energy. Layer 0 provides energy sovereignty through community-owned solar arrays and wind turbines, battery storage, smart microgrid management via open-source controllers such as OpenEMS, and peer-to-peer energy trading. Som Energia in Spain, a renewable energy cooperative with over 80,000 members, demonstrates that citizen-owned energy is commercially viable (Kunze and Becker, 2015). Brooklyn Microgrid demonstrates peer-to-peer solar energy trading at the neighborhood scale (Mengelkamp et al., 2018). Full automation requires abundant, cheap, clean energy, and the accelerationist ecological thesis holds that the answer is not degrowth but directed technological acceleration toward clean abundance (Jacobson et al., 2017). We term this vision accelerationist ecologism: the conviction that ecological sustainability is best achieved not by reducing human activity but by powering it entirely with clean, renewable, and ultimately abundant energy sources that operate within planetary boundaries (Raworth, 2017). Table˜7 summarizes the frontier energy technologies.
@@ 1 Manufacturing Energy Access to clean energy + open-source hardware designs for local fabrication
Energy is the precondition for everything. Without clean energy sovereignty, no other layer achieves long-term sustainability: manufacturing depends on power for CNC machines and 3D printers, food production depends on power for irrigation pumps and controlled-environment systems, communication depends on electricity for servers and routers, and governance depends on the infrastructure that communication provides. The gate condition for the energy layer is that a community must achieve greater than 50% renewable self-generation capacity before the energy layer can be considered operational. This threshold is not arbitrary: below 50%, the community remains structurally dependent on external energy providers and vulnerable to supply disruptions or price manipulation. Current examples demonstrate that this gate is achievable. The World Bank is accelerating renewable deployment in developing countries through its Climate Change Action Plan, with particular emphasis on distributed solar in Sub-Saharan Africa and South Asia. The European Union’s Clean Energy Package includes directives enabling energy communities to generate, store, and sell renewable energy collectively, with over 3,400 energy cooperatives now operational across Europe (Kunze and Becker, 2015). Community solar programs in the United States have grown to over 6 GW of capacity, providing clean energy access to households that cannot install rooftop panels. The Just Energy Transition Partnerships (JETPs) negotiated by the World Bank and partner governments in South Africa, Indonesia, and Vietnam (World Bank, 2023) represent the most ambitious international framework yet for financing the transition from fossil fuels to renewables in developing economies, with combined pledges exceeding \$40 billion. A significant consideration within the energy layer is nuclear waste management. Conventional approaches to nuclear waste storage, including deep geological repositories, face both technical challenges and public opposition. An unconventional but increasingly plausible solution is space disposal via reusable rockets: as launch costs decline from approximately \$10,000/kg (Space Shuttle era) to projections below \$100/kg with fully reusable vehicles such as SpaceX Starship, the economics of launching vitrified high-level waste into solar orbit or toward the Sun become tractable. We note this not as a near-term proposal but as an illustration of how technological acceleration opens solution spaces that were previously dismissed as economically absurd.
@@
In the fourth phase, the state withers as commons governance matures, and international federation handles coordination between self-governing regions. This phase is the most speculative and the most distant. It requires not only the successful operation of all Liberation Stack layers at regional scale but also the development of effective inter-regional governance protocols, the resolution of free-rider problems between federations, and the management of common-pool resources (atmosphere, oceans, biodiversity) at planetary scale. We present this phase as a theoretical horizon, not as a near-term prediction. The critical precondition for each phase is demonstrated competence at the previous level: if commons governance fails at local scale, there is no expansion to regional scale. If regional federations cannot coordinate effectively, there is no progression to inter-regional governance. This constraint is not merely prudential; it is structural. A commons governance system that has not proven itself at small scale has no basis for claiming authority at larger scale. The reversibility of each phase is essential: if a commons institution fails to deliver on its promises, the state’s functions can be restored without systemic collapse. This is the fundamental difference between progressive state dissolution and revolutionary state seizure: the former is incremental and reversible, the latter is sudden and irreversible.
The accelerationist ecological thesis is supported not only by the energy technologies already discussed but by a broader set of scientific frontiers that make the prospect of material abundance within ecological constraints increasingly credible. We survey these frontiers across four domains: econometric evidence for decoupling, biological engineering, advanced manufacturing toward post-scarcity, and complementary sciences.
@@ 6 Empirical Evidence: Commons That Scale
A framework built on care and cooperation may sound idealistic. The evidence, however, is extensive. We present six case studies spanning different domains, scales, and geographies, all demonstrating that commons-based, decentralized, cooperative systems already work at scale.
@@ 7 Roadmap, Limitations, and Conclusion
A framework without a path of implementation is merely an aspiration. Table˜12 summarizes a four-phase roadmap toward the care-centered society.
Table 12: Phased roadmap toward the care-centered society. Timelines are indicative and depend on the assumptions discussed in the text.
Phase Horizon Core milestone Key deliverables
@@
I. Seeds 2025–2030 Pilot communities Deploy complete Liberation Stacks in 100+ pilot communities; launch federated UDR pilots at municipal scale
@@
II. Roots 2030–2040 Regional infrastructure Interconnect pilots into regional commons federations; achieve energy and food sovereignty in participating regions
@@
III. Canopy 2040–2055 Full automation Robotic systems autonomously manage energy, manufacturing, agriculture, logistics; UDR operational at regional scale
@@
IV. Forest 2055+ Care-centered society Post-labor society across federations; human activity centered on care, creativity, education, art, science, community
@@
Phase I (Seeds, 2025–2030) is already underway. The technologies described in the Liberation Stack are mature, the governance models are proven, and the cost curves for renewable energy and robotic hardware are declining exponentially. The task of this phase is integration: assembling existing components into complete, self-sustaining community stacks and demonstrating that they work together. Concretely, this means deploying community solar microgrids with open-source energy management, connecting them to fablabs with robotic manufacturing capability, establishing agroecological food production augmented by autonomous farming robots, linking everything through community-owned mesh networks, governing the whole through open-source democratic platforms, and beginning to experiment with UDR at the local level: residents of pilot communities request goods and services from the productive commons rather than purchasing them on markets. The nonviolent mobilization strategy in this phase emphasizes education, fascination with AI’s potential for human liberation, and practical benefit through pilot Liberation Stacks that provide immediate material improvements. The gate condition for progression to Phase II is the successful operation of complete Liberation Stacks in at least 100 pilot communities across at least 10 countries, with measurable improvements in energy costs, food security, digital connectivity, and democratic participation relative to comparable non-participating communities. If Phase I does not achieve these metrics within the pilot communities, Phase II cannot begin. This is not arbitrary conservatism; it is the empirical validation without which scaling would be reckless.
@@
Phase II (Roots, 2030–2040) scales what works. Communities that have demonstrated viable Liberation Stacks federate into regional networks, sharing surplus energy, manufacturing capacity, knowledge, and governance tools. The key technical milestone of this phase is the deployment of autonomous robotic systems for agriculture, construction, and logistics at a scale sufficient to meet the material needs of entire regions without human labor. The key institutional milestone is the development of inter-community commons governance protocols: how do federated communities allocate shared resources, resolve conflicts, and coordinate production across scales? The nonviolent mobilization strategy in this phase shifts to noncooperation with corporate dependencies and the growth of federated commons alternatives: communities progressively withdraw from centralized utilities, corporate platforms, and extractive supply chains as commons infrastructure provides viable substitutes. Progressive state dissolution in this phase means the state creates formal legal frameworks for commons federations, recognizes inter-community governance protocols, and begins delegating operational service delivery to commons institutions that have demonstrated competence. The gate condition for progression to Phase III is energy and food sovereignty in participating regions (defined as greater than 80% of energy and food needs met by commons infrastructure), functional federated governance across at least 10 regional federations, and measurable reductions in Gini coefficient and ecological footprint relative to non-participating regions.
@@
Phase III (Canopy, 2040–2055) is the inflection point. When robotic systems can autonomously manage the entire material production chain, from energy harvesting through raw material processing, manufacturing, food production, logistics, construction, and infrastructure maintenance, the need for human participation in instrumental labor effectively vanishes. At this point, money ceases to serve its historical function as a coordination mechanism for scarce labor. Communities that have built robust commons governance can transition to full UDR: every person has access to whatever they need, produced on demand by the robotic commons, subject only to ecological sustainability constraints computed in real time by environmental monitoring systems. The nonviolent mobilization strategy in this phase emphasizes institutional consolidation: developing the governance protocols for post-monetary commons economies, negotiating resource-sharing agreements between federations, and establishing the ecological monitoring systems that enforce planetary boundaries as first-order constraints on production. Progressive state dissolution in this phase means federated commons governance handles the majority of resource allocation, and the state’s role reduces to conflict resolution between federations and external defense. The gate condition for Phase IV is the sustained autonomous operation of the complete material production chain across multiple federations, with ecological footprint demonstrably within planetary boundaries, for a period of at least five years.
Phase IV (Forest, 2055+) is the horizon. It is not a destination but an ongoing process of collective flourishing. A society in which no conscious being is compelled to perform instrumental labor is a society that must continually reinvent what it means to live well. The work of this phase is care: raising children, tending to the elderly, nurturing friendships, creating art, pursuing science, building community, exploring the cosmos, and deepening the philosophical and spiritual dimensions of human existence. This is not idleness; it is the fullest expression of human potential, liberated from the compulsion to sell one’s time in order to survive. Progressive state dissolution in this phase means the state withers as commons governance matures, and international federation handles coordination between self-governing regions.
@@
The Pareto front framework connects the roadmap to the degrowth-accelerationism debate. The phasing itself hedges against ecological risk: Phase I deploys only mature, well-characterized technologies (solar, wind, batteries), while frontier technologies (fusion, advanced geothermal, synthetic biology) enter the stack only as they demonstrate safety and sustainability. This is the degrowth wisdom: proceed cautiously with uncharacterized technologies. Simultaneously, the overall trajectory accelerates clean technology deployment: this is the accelerationist commitment. The combination is not a compromise but a rational response to a multi-objective optimization problem.
@@
This roadmap rests on several explicit assumptions that we state transparently so that the framework can be evaluated against them. First, we assume that the exponential trajectory of AI and robotic capability continues, as predicted by the universal approximation theorem and DRL convergence results and as empirically observed over the past two decades. If progress stalls or decelerates significantly, Phases III and IV will be delayed but not invalidated: the Liberation Stack provides value at every stage of automation, not only at full automation. Second, we assume that sufficient renewable energy can be deployed to power full automation. Current projections from IRENA (IRENA, 2023) and the exponential decline in solar costs support this assumption, but it requires sustained investment in energy storage, grid infrastructure, and materials recycling. Third, we assume that open-source AI remains competitive with proprietary models. The current ecosystem of open-weight models suggests that this is viable, but sustained community investment in training infrastructure is essential. Fourth, we assume that communities can develop effective commons governance at scale. The empirical evidence supports this for systems up to millions of participants, but planetary-scale federation remains an open institutional design problem. Fifth, we assume no catastrophic pandemic disrupts global supply chains during the critical scaling phases. The COVID-19 pandemic demonstrated that global supply chain disruptions can set back technological deployment by years; a pandemic of similar or greater severity during Phase II or III could significantly delay the roadmap. Sixth, we assume no major war disrupts the cooperative international environment necessary for the commons transition. The current geopolitical environment, including the conflict in Ukraine and tensions in East Asia, illustrates that great-power conflict remains a serious risk that could redirect technological investment from civilian to military applications and fragment the open-source ecosystem along geopolitical lines. Seventh, we assume that the political window for commons-based policy remains open: that democratic institutions continue to function sufficiently to enable cooperative legislation, energy community directives, and commons-friendly regulation. Democratic backsliding or authoritarian capture of state institutions would close this window. Eighth, we assume no catastrophic AI alignment failure: that frontier AI systems do not develop goals misaligned with human values in ways that resist correction. While we consider this risk low in the near term, its consequences would be severe enough to invalidate the entire framework.
@@
We acknowledge several limitations of the framework as presented with full transparency. The timeline is inherently speculative; we offer it as an indicative structure, not as a prediction. The phasing assumes a gradual, bottom-up transition through federated pilot communities, but historical transitions are rarely linear: they involve crises, reversals, and unpredictable accelerations. The framework does not fully address geopolitical resistance from existing power structures: states, corporations, and military establishments have strong incentives to prevent the decentralization of productive capacity, and the history of social movements, from the Paris Commune to the Spanish Revolution, suggests that this resistance can be lethal. The consciousness argument that justifies delegating labor to machines (Garrido-Merchán, 2024; Garrido-Merchán and Lumbreras, 2022), while philosophically rigorous, rests on the hard problem of consciousness, which remains unresolved in philosophy of mind; if future evidence were to demonstrate that certain computational architectures do give rise to phenomenal experience, the ethical calculus of machine labor would need to be revised, potentially requiring significant restructuring of the automation framework. The UDR concept requires extensive economic modeling and institutional experimentation before it can be confidently deployed at scale; we present it here as a design principle and research direction, not as a finished policy proposal. The intersectional liberation thesis, while structurally sound, requires empirical validation: measuring the actual reduction of intersectional oppression in communities that adopt commons-based infrastructure is essential for refining the framework. The preconditions framework, while an improvement over vague roadmaps, relies on threshold values (50% renewable self-generation, 80% energy sovereignty) that are necessarily approximate and may require adjustment based on pilot experience. Finally, the progressive state dissolution framework assumes a cooperative relationship between state institutions and commons movements, which may not obtain in all political contexts: authoritarian states, captured regulatory agencies, or deeply polarized polities may resist commons-building regardless of its demonstrated benefits.
@@
A world based on care, love, and the common good is not naive. What is naive is believing that a system designed for extraction will somehow produce well-being and justice. The technologies are ready. The theorems are proven. The evidence is clear. The traditions converge. What remains is the will to build. The world we describe is not behind us, in a romanticized past. It is not above us, in an impossible utopia. It is among us: in the Linux kernel, in the Wikipedia article, in the cooperative assembly, in the community solar panel, in the Mastodon server, in the shared garden. It is among us, and it is growing.

### v6

@@ 3 Bridge-Building and Intersectional Liberation
Liberalism’s greatest insight is that individual freedom matters. Anarcho-accelerationism shares this commitment without reservation. Where I diverge is on the conditions for genuine autonomy. A person who depends on a monopolistic employer for income, a monopolistic landlord for shelter, and a monopolistic platform for communication is not free in any meaningful sense (Pettit, 1997). Genuine autonomy requires material independence, which in turn requires commons-based infrastructure: community-owned solar microgrids (Farrell, 2016), self-hosted federated services that eliminate dependence on corporate platforms (Zulli et al., 2020), and platform cooperatives that eliminate dependence on extractive intermediaries (Scholz, 2016).
@@
Intersectionality, as articulated by Crenshaw (1989) and developed by Collins (2019), identifies that systems of oppression do not operate independently but intersect: race, gender, class, sexuality, disability, age, nationality, and other axes of identity produce compound disadvantages that cannot be understood by examining any single dimension in isolation. I argue that UDR, implemented through the decentralized robotic commons, constitutes the most comprehensive intersectional intervention yet proposed, because it eliminates the material basis of all axes of oppression simultaneously. UDR eliminates class by definition: when every person has direct access to goods and services from the robotic commons, the distinction between those who own the means of production and those who must sell their labor dissolves. Racial oppression loses its material leverage when employment, housing, food, healthcare, and education are provided by the commons independently of market participation. The gendered division of labor loses its material foundation when all instrumental labor is automated. When access to resources is decoupled from productive capacity, the distinction between able and disabled bodies loses its economic significance. Economic dependence on discriminatory employers is eliminated for LGBTQ+ individuals. Both the young and the elderly gain full access to the resources of the commons. And when material well-being is no longer tied to citizenship in a wealthy state, the economic desperation that drives forced migration is eliminated. Table˜6 summarizes these interventions.
@@
This progressive path embodies what I term synthetic liberty, in contrast to the negative liberty of neoliberalism. Isaiah Berlin’s classic distinction (Berlin, 1969) between negative liberty (freedom from interference) and positive liberty (freedom to act) illuminates the inadequacy of the neoliberal conception of freedom. Van Parijs (1995) argues that real freedom requires the material conditions for genuine self-determination. Synthetic liberty is the positive freedom that emerges when commons-based infrastructure provides those conditions. It is synthetic in two senses: it is constructed through deliberate institutional and technological design, and it synthesizes the insights of multiple traditions.
@@ 4 The Liberation Stack, Accelerationist Ecologism, and the Pareto Frontier
Anarcho-accelerationism proposes a concrete layered architecture for sociotechnical systems that I term the Liberation Stack, illustrated in Figure˜7. Each layer is composed of existing, mature, open-source technologies.
@@
No community is truly autonomous if it depends on external entities for energy. Layer 0 provides energy sovereignty through community-owned solar arrays and wind turbines, battery storage, smart microgrid management via open-source controllers such as OpenEMS, and peer-to-peer energy trading. Som Energia in Spain, a renewable energy cooperative with over 80,000 members, demonstrates that citizen-owned energy is commercially viable (Kunze and Becker, 2015). Brooklyn Microgrid demonstrates peer-to-peer solar energy trading at the neighborhood scale (Mengelkamp et al., 2018). Full automation requires abundant, cheap, clean energy, and the accelerationist ecological thesis holds that the answer is not degrowth but directed technological acceleration toward clean abundance (Jacobson et al., 2017). I term this vision accelerationist ecologism: the conviction that ecological sustainability is best achieved not by reducing human activity but by powering it entirely with clean, renewable, and ultimately abundant energy sources that operate within planetary boundaries (Raworth, 2017). Table˜7 summarizes the frontier energy technologies.
@@ 1 Manufacturing Energy Access to clean energy + open-source hardware designs for local fabrication
Energy is the precondition for everything. Without clean energy sovereignty, no other layer achieves long-term sustainability: manufacturing depends on power for CNC machines and 3D printers, food production depends on power for irrigation pumps and controlled-environment systems, communication depends on electricity for servers and routers, and governance depends on the infrastructure that communication provides. The gate condition for the energy layer is that a community must achieve greater than 50% renewable self-generation capacity before the energy layer can be considered operational. This threshold is not arbitrary: below 50%, the community remains structurally dependent on external energy providers and vulnerable to supply disruptions or price manipulation. Current examples demonstrate that this gate is achievable. The World Bank is accelerating renewable deployment in developing countries through its Climate Change Action Plan, with particular emphasis on distributed solar in Sub-Saharan Africa and South Asia. The European Union’s Clean Energy Package includes directives enabling energy communities to generate, store, and sell renewable energy collectively, with over 3,400 energy cooperatives now operational across Europe (Kunze and Becker, 2015). Community solar programs in the United States have grown to over 6 GW of capacity, providing clean energy access to households that cannot install rooftop panels. The Just Energy Transition Partnerships (JETPs) negotiated by the World Bank and partner governments in South Africa, Indonesia, and Vietnam (World Bank, 2023) represent the most ambitious international framework yet for financing the transition from fossil fuels to renewables in developing economies, with combined pledges exceeding \$40 billion. A significant consideration within the energy layer is nuclear waste management. Conventional approaches to nuclear waste storage, including deep geological repositories, face both technical challenges and public opposition. An unconventional but increasingly plausible solution is space disposal via reusable rockets: as launch costs decline from approximately \$10,000/kg (Space Shuttle era) to projections below \$100/kg with fully reusable vehicles such as SpaceX Starship, the economics of launching vitrified high-level waste into solar orbit or toward the Sun become tractable. I note this not as a near-term proposal but as an illustration of how technological acceleration opens solution spaces that were previously dismissed as economically absurd.
@@
In the fourth phase, the state withers as commons governance matures, and international federation handles coordination between self-governing regions. This phase is the most speculative and the most distant. It requires not only the successful operation of all Liberation Stack layers at regional scale but also the development of effective inter-regional governance protocols, the resolution of free-rider problems between federations, and the management of common-pool resources (atmosphere, oceans, biodiversity) at planetary scale. I present this phase as a theoretical horizon, not as a near-term prediction. The critical precondition for each phase is demonstrated competence at the previous level: if commons governance fails at local scale, there is no expansion to regional scale. If regional federations cannot coordinate effectively, there is no progression to inter-regional governance. This constraint is not merely prudential; it is structural. A commons governance system that has not proven itself at small scale has no basis for claiming authority at larger scale. The reversibility of each phase is essential: if a commons institution fails to deliver on its promises, the state’s functions can be restored without systemic collapse. This is the fundamental difference between progressive state dissolution and revolutionary state seizure: the former is incremental and reversible, the latter is sudden and irreversible.
The accelerationist ecological thesis is supported not only by the energy technologies already discussed but by a broader set of scientific frontiers that make the prospect of material abundance within ecological constraints increasingly credible. I survey these frontiers across four domains: econometric evidence for decoupling, biological engineering, advanced manufacturing toward post-scarcity, and complementary sciences.
@@ 6 Empirical Evidence: Commons That Scale
A framework built on care and cooperation may sound idealistic. The evidence, however, is extensive. I present six case studies spanning different domains, scales, and geographies, all demonstrating that commons-based, decentralized, cooperative systems already work at scale.
@@ 7 Roadmap, Limitations, and Conclusion
A framework without a path of implementation is merely an aspiration. Table˜12 summarizes a four-phase roadmap toward the care-centered society. Crucially, the roadmap does not specify absolute dates. The transition can only begin when a sufficient base of social support has been built through sustained education, public deliberation, and demonstrated pilot successes, so that communities choose to adopt commons-based infrastructure through informed, voluntary, democratic decision-making. This precondition is not incidental but fundamental: a peaceful transition requires that the population freely decides to undertake it, and this in turn requires a prior process of education and consciousness-raising that cannot be scheduled in advance. The durations indicated below are therefore relative estimates, contingent on the pace at which this social foundation is established.
Table 12: Phased roadmap toward the care-centered society. Durations are relative estimates from the onset of each phase; absolute timing depends on the prior achievement of sufficient social support through education and democratic deliberation.
Phase Duration Core milestone Key deliverables
@@
I. Seeds $\sim$}5 years Pilot communities Deploy complete Liberation Stacks in 100+ pilot communities; launch federated UDR pilots at municipal scale
@@
II. Roots $\sim$}10 years Regional infrastructure Interconnect pilots into regional commons federations; achieve energy and food sovereignty in participating regions
@@
III. Canopy $\sim$}15 years Full automation Robotic systems autonomously manage energy, manufacturing, agriculture, logistics; UDR operational at regional scale
@@
IV. Forest Open-ended Care-centered society Post-labor society across federations; human activity centered on care, creativity, education, art, science, community
@@
Phase I (Seeds, approximately 5 years) begins once sufficient social support has been established through education, pilot demonstrations, and democratic deliberation. The technologies described in the Liberation Stack are already mature, the governance models are proven, and the cost curves for renewable energy and robotic hardware continue to decline. The task of this phase is integration: assembling existing components into complete, self-sustaining community stacks and demonstrating that they work together. Concretely, this means deploying community solar microgrids with open-source energy management, connecting them to fablabs with robotic manufacturing capability, establishing agroecological food production augmented by autonomous farming robots, linking everything through community-owned mesh networks, governing the whole through open-source democratic platforms, and beginning to experiment with UDR at the local level: residents of pilot communities request goods and services from the productive commons rather than purchasing them on markets. The nonviolent mobilization strategy in this phase emphasizes education, fascination with AI’s potential for human liberation, and practical benefit through pilot Liberation Stacks that provide immediate material improvements. The gate condition for progression to Phase II is the successful operation of complete Liberation Stacks in at least 100 pilot communities across at least 10 countries, with measurable improvements in energy costs, food security, digital connectivity, and democratic participation relative to comparable non-participating communities. If Phase I does not achieve these metrics within the pilot communities, Phase II cannot begin. This is not arbitrary conservatism; it is the empirical validation without which scaling would be reckless.
@@
Phase II (Roots, approximately 10 years) scales what works. Communities that have demonstrated viable Liberation Stacks federate into regional networks, sharing surplus energy, manufacturing capacity, knowledge, and governance tools. The key technical milestone of this phase is the deployment of autonomous robotic systems for agriculture, construction, and logistics at a scale sufficient to meet the material needs of entire regions without human labor. The key institutional milestone is the development of inter-community commons governance protocols: how do federated communities allocate shared resources, resolve conflicts, and coordinate production across scales? The nonviolent mobilization strategy in this phase shifts to noncooperation with corporate dependencies and the growth of federated commons alternatives: communities progressively withdraw from centralized utilities, corporate platforms, and extractive supply chains as commons infrastructure provides viable substitutes. Progressive state dissolution in this phase means the state creates formal legal frameworks for commons federations, recognizes inter-community governance protocols, and begins delegating operational service delivery to commons institutions that have demonstrated competence. The gate condition for progression to Phase III is energy and food sovereignty in participating regions (defined as greater than 80% of energy and food needs met by commons infrastructure), functional federated governance across at least 10 regional federations, and measurable reductions in Gini coefficient and ecological footprint relative to non-participating regions.
@@
Phase III (Canopy, approximately 15 years) is the inflection point. When robotic systems can autonomously manage the entire material production chain, from energy harvesting through raw material processing, manufacturing, food production, logistics, construction, and infrastructure maintenance, the need for human participation in instrumental labor effectively vanishes. At this point, money ceases to serve its historical function as a coordination mechanism for scarce labor. Communities that have built robust commons governance can transition to full UDR: every person has access to whatever they need, produced on demand by the robotic commons, subject only to ecological sustainability constraints computed in real time by environmental monitoring systems. The nonviolent mobilization strategy in this phase emphasizes institutional consolidation: developing the governance protocols for post-monetary commons economies, negotiating resource-sharing agreements between federations, and establishing the ecological monitoring systems that enforce planetary boundaries as first-order constraints on production. Progressive state dissolution in this phase means federated commons governance handles the majority of resource allocation, and the state’s role reduces to conflict resolution between federations and external defense. The gate condition for Phase IV is the sustained autonomous operation of the complete material production chain across multiple federations, with ecological footprint demonstrably within planetary boundaries, for a period of at least five years.
Phase IV (Forest, open-ended) is the horizon. It represents not a fixed destination but a sustained process of institutional and cultural development. In a society where instrumental labor is no longer required, the central design challenge becomes the creation of institutions that support meaningful human activity: care work, education, scientific research, artistic creation, community governance, and the cultivation of interpersonal relationships. The reallocation of human time from instrumental to intrinsically motivated activity constitutes the substantive content of the framework’s normative vision. Progressive state dissolution in this phase means the state withers as commons governance matures, and international federation handles coordination between self-governing regions.
@@
The Pareto front framework connects the roadmap to the degrowth-accelerationism debate. The phasing itself hedges against ecological risk: Phase I deploys only mature, well-characterized technologies (solar, wind, batteries), while frontier technologies (fusion, advanced geothermal, synthetic biology) enter the stack only as they demonstrate safety and sustainability. This phasing reflects the degrowth emphasis on caution with uncharacterized technologies, while the overall trajectory reflects the accelerationist emphasis on rapid clean technology deployment. The combination constitutes a rational response to a multi-objective optimization problem rather than a compromise between the two positions.
@@
This roadmap rests on several explicit assumptions that I state transparently so that the framework can be evaluated against them. First, I assume that the exponential trajectory of AI and robotic capability continues, as predicted by the universal approximation theorem and DRL convergence results and as empirically observed over the past two decades. If progress stalls or decelerates significantly, Phases III and IV will be delayed but not invalidated: the Liberation Stack provides value at every stage of automation, not only at full automation. Second, I assume that sufficient renewable energy can be deployed to power full automation. Current projections from IRENA (IRENA, 2023) and the exponential decline in solar costs support this assumption, but it requires sustained investment in energy storage, grid infrastructure, and materials recycling. Third, I assume that open-source AI remains competitive with proprietary models. The current ecosystem of open-weight models suggests that this is viable, but sustained community investment in training infrastructure is essential. Fourth, I assume that communities can develop effective commons governance at scale. The empirical evidence supports this for systems up to millions of participants, but planetary-scale federation remains an open institutional design problem. Fifth, I assume no catastrophic pandemic disrupts global supply chains during the critical scaling phases. The COVID-19 pandemic demonstrated that global supply chain disruptions can set back technological deployment by years; a pandemic of similar or greater severity during Phase II or III could significantly delay the roadmap. Sixth, I assume no major war disrupts the cooperative international environment necessary for the commons transition. The current geopolitical environment, including the conflict in Ukraine and tensions in East Asia, illustrates that great-power conflict remains a serious risk that could redirect technological investment from civilian to military applications and fragment the open-source ecosystem along geopolitical lines. Seventh, I assume that the political window for commons-based policy remains open: that democratic institutions continue to function sufficiently to enable cooperative legislation, energy community directives, and commons-friendly regulation. Democratic backsliding or authoritarian capture of state institutions would close this window. Eighth, I assume no catastrophic AI alignment failure: that frontier AI systems do not develop goals misaligned with human values in ways that resist correction. While I consider this risk low in the near term, its consequences would be severe enough to invalidate the entire framework.
@@
I acknowledge several limitations of the framework as presented with full transparency. The timeline is inherently speculative; I offer it as an indicative structure, not as a prediction. The phasing assumes a gradual, bottom-up transition through federated pilot communities, but historical transitions are rarely linear: they involve crises, reversals, and unpredictable accelerations. The framework does not fully address geopolitical resistance from existing power structures: states, corporations, and military establishments have strong incentives to prevent the decentralization of productive capacity, and the history of social movements, from the Paris Commune to the Spanish Revolution, suggests that this resistance can be lethal. The consciousness argument that justifies delegating labor to machines (Garrido-Merchán, 2024; Garrido-Merchán and Lumbreras, 2022), while philosophically rigorous, rests on the hard problem of consciousness, which remains unresolved in philosophy of mind; if future evidence were to demonstrate that certain computational architectures do give rise to phenomenal experience, the ethical calculus of machine labor would need to be revised, potentially requiring significant restructuring of the automation framework. The UDR concept requires extensive economic modeling and institutional experimentation before it can be confidently deployed at scale; I present it here as a design principle and research direction, not as a finished policy proposal. The intersectional liberation thesis, while structurally sound, requires empirical validation: measuring the actual reduction of intersectional oppression in communities that adopt commons-based infrastructure is essential for refining the framework. The preconditions framework, while an improvement over vague roadmaps, relies on threshold values (50% renewable self-generation, 80% energy sovereignty) that are necessarily approximate and may require adjustment based on pilot experience. Finally, the progressive state dissolution framework assumes a cooperative relationship between state institutions and commons movements, which may not obtain in all political contexts: authoritarian states, captured regulatory agencies, or deeply polarized polities may resist commons-building regardless of its demonstrated benefits.
@@
In summary, this paper has introduced peaceful anarcho-accelerationism as a rigorously defined sociotechnical framework, grounded it in a 200-year intellectual tradition, proposed a concrete technical architecture with explicit preconditions and gate conditions, and provided empirical evidence that commons-based systems already operate at scale across multiple domains. The framework’s principal contribution is the integration of automation theory, commons governance, intersectional analysis, and nonviolent strategy into a coherent design philosophy with testable hypotheses and a phased implementation roadmap. Significant empirical and institutional work remains before the framework’s more ambitious claims can be validated, but the existence of large-scale commons-based systems such as Linux, Wikipedia, Mondragon, Rojava, guifi.net, and the Fediverse demonstrates that the underlying organizational principles are not merely theoretical but operational. The framework is offered as a contribution to the interdisciplinary research agenda on the governance of automation, and its hypotheses, assumptions, and limitations have been stated explicitly to facilitate critical evaluation and refinement by the scholarly community.